OpenStudy (thechristian30005):

Annie wrote the following steps to solve a pair of equations: Step 1: y = –x – 5 y = 2x + 2 Step 2: 2y = –2x –10 y = 2x + 2 Which of the following shows the correct next step to solve the equations by eliminating x?

jimthompson5910 (jim_thompson5910):

Notice how we have -2x up top and 2x down below in step 2 also notice how -2x plus 2x = (-2+2)x = 0x = 0. So effectively, adding up the x terms in step 2 will make the x terms go away.

jimthompson5910 (jim_thompson5910):

So you have do this with both sides Adding the y terms on the left side gives 2y+y = 3y adding up the constant terms on the right side gives -10+2 = -8

jimthompson5910 (jim_thompson5910):

After adding up the two equations in step 2, we are left with 3y = -8